Circular No. 9065

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2009 L3–L17, C/2009 M1, AND C/2009 M2 (SOHO) Further to IAUC 9059, additional presumed comets have been found

on SOHO website images, all being Kreutz sungrazers except for C/2009 L8 and C/2009 L13 (Kracht group). C/2009 L3 was very diffuse and slightly elongated (mag ∼ 7.5). C/2009 L4 was stellar in appearance (mag ∼ 6.5) in C3 images but quite large and very condensed with no visible tail in C2 images; COR2-A STEREO images show a bright starlike head and a very faint, thin tail. C/2009 L5, C/2009 L7, C/2009 L11 (also described as small), and C/2009 L15 were slightly diffuse (peak magnitudes ∼ 7.5, 7.5–8, 8, and 8, respectively). C/2009 L6 was stellar in appearance (mag ∼ 6.5–7) in C3 images, and condensed with no tail in C2 images. C/2009 L8 (mag ∼ 7–7.5) and C/2009 L13 (mag ∼ 7.5–8) were small and stellar in appearance. C/2009 L9 and C/2009 L12 were very diffuse (mag ∼ 8). C/2009 L10 (slightly diffuse), C/2009 L14 (quite diffuse), and C/2009 M1 (quite diffuse) were very faint (mag ∼ 8.5). C/2009 L16 was quite diffuse (mag ∼ 7.5). C/2009 L17 was stellar in appearance (mag ∼ 6.5) in C3 images, but quite diffuse with a short, faint tail in C2 images. C/2009 M2 was small and stellar in appearance (mag ∼ 7) in C3 images, and showed a hint of a very faint tail in C2 images.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 L3 June 3.785 4^{h}45.^{m}2 +20^{o}33′ C2  RK 2009-O19
C/2009 L4  3.888 4 48.7 +18 35 C3/2* BZ 2009-O19
C/2009 L5 4.893 4 49.3 +20 39 C2 RK 2009-O19
C/2009 L6  5.988 4 55.2 +19 25 C3/2 BZ 2009-O19
C/2009 L7  1.381 4 35.2 +20 12 C2  BZ 2009-O21
C/2009 L8  6.979 4 51.8 +22 41 C2  TH 2009-O21
C/2009 L9  7.229 4 56.6 +20 51 C2  GS 2009-O21
C/2009 L10  7.329 4 58.4 +20 57 C2  BZ 2009-O21
C/2009 L11  7.671 5 00.3 +20 56 C2  MK 2009-O21
C/2009 L12  8.893 5 02.7 +20 54 C2  MU 2009-O22
C/2009 L13  9.185 5 12.1 +23 26 C2  RK 2009-O22
C/2009 L14  10.185 5 09.8 +21 09 C2  BZ 2009-O22
C/2009 L15  10.910 5 12.2 +21 11 C2  RM 2009-O22
C/2009 L16  14.393 5 25.8 +21 22 C2  MK 2009-P01
C/2009 L17  15.013 5 25.0 +20 18 C3/2 MU 2009-P01
C/2009 M1 21.838 5 55.1 +21 42 C2 BZ 2009-P01
C/2009 M2  22.238 5 56.5 +20 42 C3/2 MK 2009-P01
2009 August 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green